# Maria Bazotte, Margarida Carvalho, and Thibaut Vidal

### v3

@@ 6.2 Comparative results for NDFPP-Selection
0010101010202020203030303040404040505050506060606070707070808080809090909010010010010000101010102020202030303030404040405050505060606060707070708080808090909090100100100100Instances (%)SAA Transformed program $\left(\overline{gap}(\overline{x})(\%)\right)$DEP Original program $\left(GAP(\%)\right)$
Comparison of the optimality gap obtained by the SAA for the transformed program and the DEP of the original program.
Table 6.2 shows the experimental results of the 4- and 5-facility instances. In the table, $|K|$ and $|K_{t}|$ are the set of scenarios in the DEP of the original program (Program (2)) and in the DEP of our transformed Program (6), respectively. Moreover, $GAP(\%)$ is the optimality gap for the DEP. For the DEP, all runs reached the time limit of 2 hours. We followed the procedure explained in Section 3.1 for the SAA to obtain the feasible solution $\overline{x}$ and its estimators (see Appendix 8 for details). Even though the number of scenarios $|K_{t}|$ of the transformed program is extremely large (i.e.,$|\mathcal{D}|{|\mathcal{W}|}^{|\mathcal{N_{F}}||\mathcal{P}|}$), the number $|K|$ of scenarios of the DEP is small (i.e., $|\mathcal{D}|{|\mathcal{W}|}^{|\mathcal{N_{F}}|}$) for the considered instances. Thus, we can compute the real stochastic value of the feasible solution $\overline{x}$, i.e., $v(\overline{x})=c^{\top}\overline{x}+g(\overline{x})$. For the SAA, we solve $M=50$ problems with $N=750$ sample scenarios because these values sufficed to obtain estimators with low variance and feasible, high-quality solutions. We present the over-estimator of the relative optimality gap, i.e., $\overline{gap}(\overline{x})(\%)={100\%\times[v(\overline{x})-\overline{v}_{M}^{N}]}/{v(\overline{x})}$. Because the lower-bound estimator $\overline{v}_{M}^{N}$ is a random variable, this over-estimator of the optimality gap may assume negative values (Verweij et al. 2003). The computational time of the SAA is in seconds. For all runs, the SAA finished before the time limit of 2 hours. For comparison, we display the gap between the upper-bound values obtained by the proposed method and the DEP, i.e., $GAP_{UB}(\%)={100\%\times[UB-v(\overline{x})]}/{UB}$, and the gap between their lower bounds, i.e., $GAP_{LB}(\%)={100\%\times\left(LB-\overline{v}_{M}^{N}\right)}/{LB}$, where $UB$ and $LB$ are the upper and lower bounds for the DEP, respectively. We summarize our main findings in Results 1 and 2.
Result 1
The proposed method produces a feasible solution $\overline{x}$ with a considerably low over-estimator of the optimality gap, $\overline{gap}(\overline{x})(\%)$ is between $0\%$ and $1\%$ for all instances. The lower bound $\overline{v}_{M}^{N}$ and the optimality gap estimators have low variability, as shown by the standard deviation $\hat{\sigma}_{N,M}$ of the lower-bound estimator, which is inferior to $0.5\%$ of $\overline{v}_{M}^{N}$ for all instances (see Appendix 16 for details). For the considered instances, the computational time required by the proposed method was inferior to 2000s in 87% of the cases, with an average of around 990s.
Result 2
The proposed method’s over-estimator of the optimality gap is significantly smaller than the DEP’s optimality gap (the difference is around $24\%$ on average). The proposed method obtained these optimality gaps in significantly less time than the DEP, and, within the time limit, consistently found feasible solutions with slightly better values than those from the DEP.
Result 1 highlights the good performance of the proposed method. Result 2 shows that it outperformed the DEP. Figure 6.2, which shows the optimality gap for the DEP and the proposed approach, corroborates this analysis. The proposed method obtained optimality gaps close to $0\%$ for all instances, whereas the DEP obtained optimality gaps inferior to $10\%$ only for $22\%$ of the instances. The DEP obtained optimality gaps of $10\%$$20\%$ for around $56\%$ of the instances, of $20\%$$30\%$ for around $8\%$ of the instances, and of $100\%$ for around $14\%$ of the instances.
As previously stated, the transformed Program (6) has more scenarios than the DEP of the original program. Nonetheless, we apply the SAA to our transformed program to efficiently solve different problems with fewer scenarios. The proposed approach obtained high-quality solutions by sampling only $N=750$ scenarios for each of the $M=50$ SAA problems, which, in most cases, is less than the number $|K|$ of scenarios of the DEP. Despite the number of scenarios, the proposed method solves mathematical programs smaller than those obtained with the DEP (in terms of their number of variables and constraints). The DEP’s MILP formulation contains numerous big-M constraints due to the linearization of the probability function (4), which leads to weak linear relaxations. In contrast, the transformation function applied to obtain the NDFPP-Selection formulation (see Sections 4.1 and 5.1.1) includes no extra variables or constraints.
6.3 Analysis for NDFPP-Binomial, NDFPP-Discrete, and NDFPP-Normal
In this section, we analyze the statistical estimators and feasible solution $\overline{x}$ of the SAA applied to our transformed program for NDFPP-Binomial, NDFPP-Discrete, and NDFPP-Normal. In these cases, the DEP of the original program (Program (2) for discrete distributions) is nonlinear nonconvex, and the linearization of Bhuiyan et al. (2020) cannot be applied. Moreover, note that our transformed Program (6) for these variants has continuous random variables, making the explicit enumeration of scenarios impossible and requiring the implementation of sampling techniques.
The DEP of the original program has a small number $|K|$ of scenarios for the NDFPP-Binomial and NDFPP-Discrete instances. Thus, we compute $v(\overline{x})$ for these cases. The number of scenarios of the DEP of the original program is infinite for NDFPP-Normal because the endogenous random variables are continuous. Consequently, we compute the estimator $\hat{v}_{N{{}^{\prime}}}(\overline{x})$ of $v(\overline{x})$. Table 6.3 lists the experimental results of the 4- and 5-facility instances. The over-estimator of the optimality gap $\overline{gap}(\overline{x})(\%)$ is defined as in the previous section for NDFPP-Binomial and NDFPP-Discrete. For NDFPP-Normal, $\overline{gap}(\overline{x})(\%)={100\%\times[\hat{v}_{N{{}^{\prime}}}(\overline{x})-\overline{v}_{M}^{N}]}/{\hat{v}_{N{{}^{\prime}}}(\overline{x})}$. We set $M=50$ and $N=750$ for the same reason as in the previous section. For NDFPP-Normal, we use $N{{}^{\prime}}=5\times 10^{4}$. The SAA finished for all instances before reaching 3 hours. The result is as follows:
Result 3
The proposed method obtained a feasible solution $\overline{x}$ with a significantly low statistical over-estimator of the optimality gap $\overline{gap}(\overline{x})(\%)$ between $-1\%$ and $1\%$ for all instances of the problem versions. The SAA’s estimators have high confidence, as shown by the standard deviation $\hat{\sigma}_{N,M}$ of the lower-bound estimator, which is inferior to $0.5\%$ of $\overline{v}_{M}^{N}$ for all cases, and the standard deviation $\hat{\sigma}_{N{{}^{\prime}}}(\overline{x})$ of the upper bound estimator, which is inferior to $0.2\%$ of $\hat{v}_{N{{}^{\prime}}}(\overline{x})$ for all NDFPP-Normal instances (see Appendix 17 for details). The computational time was, on average, around 1500s for NDFPP-Binomial, 3000s for NDFPP-Discrete, and 1200s for NDFPP-Normal.
[…]
@@ 6.3 Analysis for NDFPP-Binomial, NDFPP-Discrete, and NDFPP-Normal
The proposed method found high-quality solutions and estimators for the considered problem variants. The computational time of this approach for the NDFPP-Binomial, NDFPP-Discrete, and NDFPP-Normal variants exceeded that for the NDFPP-Selection variant. Solving the NDFPP-Discrete variant required the longest computational time. The transformation function used in this case (i.e., inverse transformation for discrete random variables; see Section 4.2.1) includes numerous binary variables and big-M constraints in the formulation. The initial formulation of NDFPP-Discrete contained bilinear terms, which were linearized by adding more big-M constraints in the formulation. For the same reason, the computational time of the proposed method for NDFPP-Binomial exceeded that for NDFPP-Selection. The computational time of the proposed method for solving NDFPP-Discrete increases with the number of capacity levels (possible realizations of the random variables). Therefore, a trade-off exists between how sample size affects the performance and confidence of the SAA’s statistical estimators and the associated computational time. For NDFPP-Normal, the number of evaluation scenarios $N{{}^{\prime}}$ is considerably large, which increases the computational time of the proposed method. Moreover, the standard deviation of the optimality gap is slightly larger than for the remaining versions because it includes the variability of $\hat{v}_{N{{}^{\prime}}}(\overline{x})$.
6.4 Value of the Stochastic Solution
@@
This section analyzes the relative VSS of all problem variants. The relative VSS is defined as ${(\textrm{EEV}-v^{*})}/{\textrm{EEV}}$, where $v^{*}$ is the optimal value of the recourse problem, and the EEV is the expected result of using the EV solution. Thus, $\textrm{EEV}=v(\hat{x}^{*})=c^{\top}\hat{x}^{*}+g(\hat{x}^{*})$, where $\hat{x}^{*}$ is the optimal solution of the EV. In Appendix 13, we detail the EVs (i.e., the deterministic problems considering the average case scenario) according to the problem variants. We use the feasible solution $\overline{x}$ obtained by the SAA applied to our transformed program to compute under estimators of the VSS. We have the under estimator $\textrm{VSS}_{1}={[\textrm{EEV}-v(\overline{x})]}/{\textrm{EEV}}$ for NDFPP-Selection, NDFPP-Binomial, and NDFPP-Discrete, for which we can enumerate the DEP scenarios. For NDFPP-Normal, we compute $\textrm{VSS}_{2}={[\overline{\textrm{EEV}}-\hat{v}_{N{{}^{\prime}}}(\overline{x})]}/{\overline{\textrm{EEV}}}$, where $\overline{\textrm{EEV}}=\hat{v}_{N{{}^{\prime}}}(\hat{x}^{*})$ and $N{{}^{\prime}}=5\times 10^{4}$. Table 6.4 summarizes the under estimators of the VSS for the 4- and 5-facility instances.
[…]
@@
The proposed method obtained better solutions than the EV in all cases. Thus, solving the recourse problem instead of the EV is also advantageous in the context of endogenous uncertainty, where the distribution defined by $\hat{x}^{*}$ is not the same as that defined by $\overline{x}$ or $x^{*}$. Different instances of NDFPP-Selection showed a high VSS, around $30\%$ or more. Some instances of NDFPP-Binomial and NDFPP-Discrete also presented high VSSs, whereas others had VSSs of $1\%$$4\%$. Moreover, although the VSS for NDFPP-Normal approached $0\%$ for the 5-facility instances, it hovered around $3\%$$6\%$ for the 4-facility instances. Although the VSS is primarily instance-dependent, it is also affected by the problem (e.g., NDFPP-Selection and NDFPP-Normal for 15-node instances).
@@ 7 Conclusions
We proposed a general method for addressing 2-SMIPD that uses the random variable transformation. Our methodology (i) works for a general class of programs with endogenous uncertainty; (ii) obtains a stochastic program with exogenous uncertainty for which solution procedures are well-studied; (iii) yields MILPs or MICPs with exogenous uncertainty in many cases; (iv) works for both discrete and continuous endogenous random variables; and (v) is applicable to problems with independent and correlated random variables. We investigated techniques to define appropriate transformations of random variables to treat cases where the decision-dependent random variables follow classical probability distributions. We showed the existence of a simple transformation for the case of the discrete selection of distributions, which is the most commonly studied in previous research. We also demonstrated the proposed methodology by applying it to the NDFPP, a broadly studied problem in the literature. In the computational experiments, the SAA applied to our transformed program produced feasible, high-quality solutions and high-confidence estimators in a short computational time.
Numerous research perspectives stem from this work. Future work could focus on (i) exploring additional techniques to define transformation functions; (ii) investigating valid inequalities for the discrete random variables inverse transformation; (iii) analyzing how endogenous uncertainty affects the definition of the VSS; (iv) examining stochastic programs with endogenous random variables for which the probability distribution is not known a priori; and (v) extending the NDFPP to account for assumptions commonly investigated in previous research, such as post-disruption edge failure, protection (Viswanath et al. 2004, Flach 2010, Peeta et al. 2010, Prestwich et al. 2014, Du and Peeta 2014), and facility location (Li et al. 2021, 2022).

### v4

@@ 5.2 Experimental Results
The SAA-TP found high-quality solutions and estimators for the considered problems. The computational time of this approach for the NDFPP-Binomial, NDFPP-Discrete, and NDFPP-Normal exceeded that for the NDFPP-Selection. Solving the NDFPP-Discrete required the longest computational time. The transformation used in this case (i.e., inversion for discrete random variables; see Section 4.1.1) includes numerous binary variables and big-M constraints in the formulation. The initial formulation of NDFPP-Discrete contained bilinear terms, which were linearized by adding more big-M constraints in the formulation. For the same reason, the computational time of the SAA-TP for NDFPP-Binomial exceeded that for NDFPP-Selection. The computational time of the SAA-TP for solving NDFPP-Discrete increases with the number of capacity levels (possible realizations of the random variables). Therefore, a trade-off exists between how sample size affects the performance and confidence of the SAA’s statistical estimators and the associated computational time. For NDFPP-Normal, the number of evaluation scenarios $N{{}^{\prime}}$ is considerably large, which increases the computational time of the SAA-TP. Moreover, the standard deviation of the optimality gap is slightly larger than for the remaining versions because it includes the variability of $\hat{v}_{N{{}^{\prime}}}(\overline{x})$.
5.2.4 Value of the Stochastic Solution.
@@
This section analyzes the relative VSS of all problem variants. The relative VSS is defined as ${(\textrm{{{{EEV}}}}-v^{*})}/{\textrm{{{{EEV}}}}}$, where $v^{*}$ is the optimal value of the recourse problem, and EEV is ${v(\hat{x}^{*})=c^{\top}\hat{x}^{*}+g(\hat{x}^{*})}$, with $\hat{x}^{*}$ the optimal solution of the EV. In Appendix 14, we detail the EVs (i.e., the deterministic problems with the average case scenario) according to the problem variants. We use the feasible solution $\overline{x}$ obtained by the SAA-TP to compute under estimators of the VSS, and we detail this computation in Appendix 20. We resume our main findings in Result 4.
Result 4
@@
The SAA-TP obtained better solutions than the EV in all cases. Thus, solving the recourse problem instead of the EV is also advantageous in the context of endogenous uncertainty, where the distribution defined by $\hat{x}^{*}$ is not the same as that defined by $\overline{x}$ or $x^{*}$. Different instances of NDFPP-Selection showed a high VSS, around $30\%$ or more. Some instances of NDFPP-Binomial and NDFPP-Discrete also presented high VSSs, whereas others had VSSs of $1\%$$4\%$. Although the VSS for NDFPP-Normal approached $0\%$ for the 5-facility instances, it hovered around $3\%$$6\%$ for the 4-facility instances. Although the VSS is primarily instance-dependent, it is also affected by the problem (e.g., NDFPP-Selection and NDFPP-Normal for 15-node instances have considerably different VSSs). Finally, Appendix 20 contains the complete results.
@@ 6 Conclusions
We proposed a general method for addressing 2-SMIPD that uses random variable transformation. Our methodology (i) works for a general class of programs with endogenous uncertainty; (ii) obtains a stochastic program with exogenous uncertainty for which solution procedures are well-studied; (iii) yields MILPs or MICPs with exogenous uncertainty in many cases; (iv) works for both discrete and continuous endogenous random variables; and (v) is applicable to problems with independent and correlated random variables. We investigated techniques to define appropriate transformations of random variables to treat cases where the decision-dependent random variables follow classical probability distributions. We showed the existence of a simple transformation for the case of the discrete selection of distributions, which is the most commonly studied in previous research. We also demonstrated the proposed methodology by applying it to the NDFPP, a broadly studied problem. In the computational experiments, the SAA-TP produced feasible, high-quality solutions and high-confidence estimators in a short computational time.
Numerous research perspectives stem from this work. Future work could focus on (i) exploring additional techniques to define transformation functions; (ii) investigating approaches to deal with the current limitations of the method when $q(\omega)$, $W(\omega)$ and/or $H(\omega)$ are random, such as exploring valid inequalities for the discrete random variables inverse transformation; (iii) analyzing how endogenous uncertainty affects the definition of the VSS; (iv) examining stochastic programs with endogenous random variables for which the probability distribution is not known a priori; and (v) extending the NDFPP to account for assumptions investigated in previous research, such as post-disruption edge failure, protection (see, e.g., (Du and Peeta 2014)), and facility location (Li et al. 2021, 2022).